x+4y=5 how would you graph this
Answer by kiljoy(51)

You need to start plugging in values for x or y and solving the equation.
x+4y=5
x=1
4y=5-1
y=1
This gives you a point at (1,1)
x=-3
-3+4y=5
4y=8
y=2
Now you have a point at (-3,2)
Plot those points and draw a line through it.